OpenStudy (anonymous):

HELP!GIVING OUT MEDALS! 1.If a 40 kg person is accelerating at 10 m/s2, then the force acting on him or her is 400 N A. true or B.false 2.How much force is needed to make a 60 kg object accelerate at a rate of 2.0 m/s2? A. 62 N B.62 kg C.120 N D.120 kg 3.What is the acceleration of a 30 kg object when 180 N of force is applied to it? A.6.0 m/s2 B.60 m/s2 C.210 m/s2 D.5,400 m/s2 4.Which of the following equations show how you would solve for mass if you knew the force acting on an object & the acceleration of an object? A.m=f/a B.m= fa C.m=a/f D.m=f-a

OpenStudy (gudden):

in the first one...as F = MA so yes force is 400 N

OpenStudy (gudden):

answer of second is C) 120N

OpenStudy (gudden):

and since F =MA therefore A = F/M so answer of third would be 180/30 = 18/3 = 6 i.e A

OpenStudy (gudden):

since F = MA therefore m = f/a hence answer is a)... for the 4th question..

OpenStudy (gudden):

@lequestions... if u have any problem with understanding any of these u could ask it..

OpenStudy (anonymous):

thank you :) @gudden

OpenStudy (gudden):

do u have any problem in understanding any of these ??

OpenStudy (gudden):

@lequestions .... if u don't have any question...left... then my... medal... :p